Country A can produce 80 of good X and zero of Y, or 70 of good Y and zero of X (or any linear combination between). Country B can produce 140 of good X and zero of Y, or 100 of good Y and zero of good X (or linear combination). Country A has a comparative advantage in good __ and would trade if the international price is different from __units of X for one unit of Y.
Answer #1

Answer)

(i) Good Y

(ii) 1.14 units of good X for 1 unit of good Y

The production pattern of both countries can be represented in the tabular format as follows:

Country A

Good X Good Y
80 70

Country B

Good X Good Y
140 100

Comparative advantage is the advantage that a country possesses when it can produce a good at less cost as compared to another country. For this, we calculate the opportunity cost of producing each good in Country A

The opportunity cost of producing good X = (70/80) = 0.875 units pf good Y

The opportunity cost of producing good Y = (80/70) = 1.14 units of good X

Now, we calculate the opportunity cost of producing each good in Country B

The opportunity cost of producing good X = (100/140) = 0.71 units pf good Y

The opportunity cost of producing good Y = (140/100) = 1.4 units of good X

As we can see that the opportunity cost of producing good Y is low in Country A as compared to country B.So, Country A has comparative advantage in the production of good Y.

It would trade only if it could get more than the opportunity cost.For example, if country B is willing to give 1.2 units of good X to country A in return for 1 unit of good Y, then Country A would trade.